OpenStudy (anonymous):

average movie prices in the United states are in general, lower than in other countries. It would cost $78.60 to buy three tickets in Japan plus two tickets in Switzerland. Three tickets in Switzerland plus two tickets in Japan would cost $74.45 How much does an average movie ticket cost in each of these countries?

OpenStudy (anonymous):

let cost of one ticket in Japan = x $ let cost of one ticket in Switzerland = y $ then 3x + 2y = 78.60 -------------(1) 2x + 3y = 74.45 -------------(2) Multiply (1) by 2 and (2) by -3 and then add the resulting equations 6x + 4y = 157.20 -6x - 9y = -223.35 ------------------ - 5y = -66.15 y = -66.15/-5 y = 13.23 substitute y = 13.23 in (1) 3x + 2(13.23) = 78.60 3x = 78.60 - 26.46 x = 52.14/3 x = 17.38

OpenStudy (anonymous):

so cost of one ticket in Japan = x $ = $ 17.38 and cost of one ticket in Switzerland = y $ = $ 13.23